### v2

# Transformation equation for frames undergoing non-uniform acceleration such as SHM and rotational motion

Abstract
The Lorentz transformation equations provide us a set of relations between the spacetime coordinates as observed from two different inertial frames. In case one of the frames is moving with a uniform rectilinear acceleration, we have Rindler’s transformation equations under such a situation. In the present work, we extend the Rindler’s equations to a situation where we generally have non-uniform acceleration. After that, we consider the non-inertial frame to undergo simple harmonic motion (SHM), and as a second case, we consider the non-inertial frame to move uniformly along a circle. This set of transformation equations will have applications in various branches of Physics and in general in Astrophysics.
Keywords: Relativity; Frame of references; Rindler transformation; Non-uniform motions; Circular motions.
1 Introduction
In special theory of relativity, the coordinate transformation relations for a frame of reference moving with uniform velocity with respect to a rest frame are known as the Lorentz transformation [1]. These relations play a vital role during the investigation of a test particle in a frame moving with a uniform velocity. In the case of an accelerated frame, the general transformation relations are presented by Nelson [2, 3]. The coordinate transformations for uniformly accelerated frame with constant acceleration is known as the Rindler transformation, and the consequences of such a coordinate frame have been studied by different researchers [4, 5, 6, 7, 8, 9]. In the Rindler transformation, the initial velocity is considered as zero. However, the coordinate transformation for an accelerated frame with non-zero initial velocity has been discussed by Yi [6]. Additionally, Brahma and Sen [10] analyzed the Rindler transformation having acceleration along an arbitrary direction on the XY plane. In the present work, we investigate the coordinate transformation relations for a non-uniform acceleration such as a frame executing simple harmonic motion (SHM) and subsequently the transformation relations for a frame having uniform circular motion (UCM). However, in special relativity, when the motion of a test particle is subject to a restoring force along the $X$-direction, the mass of the particle can not be considered as constant. But it changes according to the law of special theory of relativity, to which MacColl [11] calculated a general expression of the time period of the relativistic oscillation, which is a function of the total energy of the system. Such a relativistic harmonic oscillation can be used to admit a consistent procedure for the quantization of the system [12]. However, the formal investigation on the harmonic oscillator in relativity using the Lagrange formalism [13] and the Hamiltonian formalism [14] indicates that even though the term “harmonic oscillator” has been generally used, but the results obtained during the calculations are no longer harmonic in nature. Instead, it is “anharmonic” in nature, which is due to the relativistic effect on mass and energy. It indicates that if the maximum velocity of the SHM is comparable to the speed of light, then the harmonic oscillation becomes an anharmonic in nature. The nature of the relativistic harmonic oscillation and the associated anharmonicity due to the relativistic effect have been experimentally studied using ultracold lithium atoms in the third band of an optical lattice and reported by Fujiwara et al. [15]. Here, it has been stated that the measured worldline shapes, relativistic anharmonicity, monopole oscillations, and relativistic dephasing of oscillator ensembles are in good agreement with the theoretical predictions of the relativistic harmonic oscillation. It is to be mentioned that, although the simple harmonic oscillation has been studied extensively in the regime of relativity but, only a few works on relativistic coordinate transformation between an inertial frame and a frame executing SHM have been reported in literature, to our knowledge. For example, Perepelkin et al. [16] discussed the transformation relation for a frame having velocity which is a sinusoidal function of time by using the generalized Galilean transformation. Therefore, it would be very much important and appropriate, to study the transformation relations for the frame undergoing SHM through a more general procedure. This motivated us to calculate such coordinate transformation relations in the present work.
So in the present context, in order to carry out the calculations to obtain the coordinate transformations between an inertial frame and a frame executing SHM, we have to first derive Rindler like transformation equations for coordinates frames having non-uniform acceleration. Since the circular motion of an object can be described by two SHMs with same amplitudes in perpendicular directions having phase difference of $\pi/2$ radians so, the transformation relations for the frame executing SHM can be used to understand the dynamics of a moving frame in a circular path. As a result, we obtain the coordinate transformation relations for a frame of reference moving in a circular path. The results in the present work will have immense applications in the study of the test particles in non-uniformly accelerated frames such as frames executing SHM and the frames moving in a circular path.
We organized our paper as: In Section 2 we obtain the transformation relations for a frame of reference executing simple harmonic motion (SHM) along the $X$-direction. Additionally, we also calculated the similar transformation relations for the frame of reference executing SHM along the $Y$-direction with initial phase of $\pi/2$ radians. And then in Section 3, as mentioned above since the circular motion can be described by using the two SHMs in perpendicular direction and phase difference of $\pi/2$ radians, towards the end we obtained the transformation relations for a frame that is moving in a circular path. In Section 4 we compared our results with the existing coordinate transformation relations for the frames executing SHM and UCM available in literature. At last, in Section 5 we discussed the results obtained in the above calculations and concluded our findings.
2 Transformation relations for a frame of reference executing SHM
The original Rindler transformation equation is meant for constant acceleration. However, in the present work, we shall derive the transformation relations in which proper acceleration or coordinate acceleration is a function of time. Most specifically for a frame executing simple harmonic motion (SHM) the coordinate acceleration is a sinusoidal function of time. Below, we derive the transformation relations for the case when the frame is oscillating along the $X$-direction having displacement proportional to some sinusoidal function of angular frequency $\omega$.
The two inertial frames $S$ and $S^{\prime}$ will have the transformation law for acceleration as [8]:
$a^{\prime}_{x}=\frac{\left(1-\frac{v^{2}}{c^{2}}\right)^{3/2}}{\left(1-\frac{u_{x}\,v}{c^{2}}\right)^{3}}a_{x}$ (1)
where $u_{x}=\frac{dx}{dt}$, $a_{x}=\frac{du_{x}}{dt}$ and $a^{\prime}_{x}=\frac{du^{\prime}_{x}}{dt^{\prime}}$. Therefore, for a particle that has acceleration $a^{\prime}_{x}$ with respect to an inertial frame $S^{\prime}$ which instantaneously accompanies the particle $(u_{x}=v)$, we can write:
$\frac{a_{x}}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)^{3/2}}=a^{\prime}_{x}$ (2)
In case of Castillo & Sanchez [8] they assumed that the acceleration $a^{\prime}_{x}$ is constant and obtain the Rindler’s transformation equation [17]. However, in our present case we shall consider $a^{\prime}_{x}$ to be a function of $(x^{\prime},t^{\prime})$ and proceed as follows.
From equation (2) we can write,
$\frac{du_{x}/dt}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)^{3/2}}=a^{\prime}_{x}$
or $\frac{du_{x}}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)}=a^{\prime}_{x}\,\left(1-\frac{u_{x}^{2}}{c^{2}}\right)^{1/2}dt$ (3)
The quantity $u_{x}$ in the right hand side is actually representing $v$. So we can identify $\left(1-\frac{u_{x}^{2}}{c^{2}}\right)^{1/2}dt=d\tau$, where $\tau$ is the proper time. Also, as mentioned by Castillo & Sanchez [8] the time $t^{\prime}$ as measured in $S^{\prime}$ frame is actually the proper time $\tau$. So we can finally write:
$\frac{du_{x}(t,x)}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)}=a_{x}^{\prime}(\tau,x^{\prime})d\tau$ (4)
Integrating on both side,
$\int\frac{du_{x}(t,x)}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)}=\int a_{x}^{\prime}(\tau,x^{\prime})d\tau$
or $c\tanh^{-1}\left(\frac{u_{x}}{c}\right)=\int a_{x}^{\prime}(\tau,x^{\prime})d\tau$
or $u_{x}=c\tanh\Big[\frac{1}{c}\int a_{x}^{\prime}(\tau,x^{\prime})d\tau\Big]$ (5)
Further, from equation (5) we can obtain as:
$dx$ $=c\tanh\Big[\frac{1}{c}\int a_{x}^{\prime}(\tau,x^{\prime})d\tau\Big]\,dt$
$\text{or\,\,\,\,\,\,\,}x$ $=\int c\tanh\Big[\frac{1}{c}\int a_{x}^{\prime}(\tau,x^{\prime})d\tau\Big]\,dt$ (6)
Also we wrote earlier $d\tau=\sqrt{1-\frac{u_{x}^{2}}{c^{2}}}dt$ and the expression for $u_{x}$ is available in equation (5) we can write :
$t=\int\Big[1-\frac{1}{c^{2}}\Big\{c\tanh\Big[\frac{1}{c}\int a_{x}^{\prime}(\tau,x^{\prime})d\tau\Big]\Big\}^{2}\Big]^{-1/2}d\tau$ (7)
Introducing dummy variables $\tau_{1}$ and $\tau_{2}$, in general we can write :
$t$ $=\int_{0}^{\tau}\Big[1-\frac{1}{c^{2}}\Big\{c\tanh\Big[\frac{1}{c}\int_{0}^{\tau_{1}}a_{x}^{\prime}(\tau_{2},x^{\prime})d\tau_{2}\Big]\Big\}^{2}\Big]^{-1/2}d\tau_{1}$ (8)
Equations (6) and (7) with appropriate constants of integration will help us to find appropriate co-ordinate transformation equations where the acceleration is not constant, but varying with time.
However, in our present case, we take $a^{\prime}_{x}$ to be a function of timetime and we shall consider two standard cases: (i) the frame $S^{\prime}$ undergoing Simple Harmonic Motion (SHM) and (ii) the frame $S^{\prime}$ under going Circular Motion.
2.1 The frame $S^{\prime}$ undergoing SHM in $X$-direction:
In the first case, we consider that acceleration $a^{\prime}_{x}$ is a function of proper time $\tau$ or $(t^{\prime})$ and represents a simple harmonic motion (SHM). So we may take a displacement $x^{\prime}=r_{0}\sin(\omega\tau)$ such that the velocity of the oscillation $u^{\prime}_{x}=\frac{dx^{\prime}}{d\tau}=r_{0}\omega\cos(\omega\tau)$ and the acceleration $a^{\prime}_{x}=\frac{d^{2}x^{\prime}}{d\tau^{2}}=-r_{0}\omega^{2}\sin(\omega\tau)$, where $r_{0}$ is the amplitude of the SHM and $\omega$ is the angular frequency. Therefore, substituting $a_{x}^{\prime}=-r_{0}\omega^{2}\sin(\omega\tau)$ in equation (4) we obtain:
$\frac{du_{x}}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)}=-r_{0}\omega^{2}\sin(\omega\tau)\,d\tau$ (9)
Now integrating both side of the equation (9) we obtain as:
$\int\frac{du_{x}}{\left(1-\frac{u_{x}^{2}}{c^{2}}\right)}=-r_{0}\omega^{2}\int\sin(\omega\tau)\,d\tau$
or $c\,\tanh^{-1}\left(\frac{u_{x}}{c}\right)=r_{0}\omega\cos(\omega\tau)+K_{1}$
or $u_{x}=c\,\tanh\left[\frac{r_{0}\omega}{c}\cos(\omega\tau)+\frac{K_{1}}{c}\right]$ (10)
where $K_{1}$ is an integration constant. Now, we consider that the displacement vector (of the $S^{\prime}$ frame) is purely oscillatory in nature, so the three velocity of the $S^{\prime}$ frame with respect to the inertial observer ($S$ frame) should be also oscillatory in nature with zero base line. Actually one can show $\tanh{\left[\frac{r_{0}\omega}{c}\cos(\omega\tau)+\frac{K_{1}}{c}\right]}$ function is oscillatory in nature with baseline which depends upon our choice of the constant $K_{1}$. But if we set $K_{1}=0$, then the baseline also vanishes to zero. So we can safely assume $K_{1}$ to be equal to 0.
Therefore, from equation (10) putting $K_{1}=0$ we can write as follows:
$u_{x}$ $=\frac{dx}{dt}=c\,\tanh\left[\frac{r_{0}\omega}{c}\cos(\omega\tau)\right]$ (11)
$\text{or\,\,\,\,\,}dx$ $=c\,\tanh\left[\frac{r_{0}\omega}{c}\cos(\omega\tau)\right]dt$ (12)
So, similarly as before, using the relation : $dt=\Big(1-\frac{u_{x}^{2}}{c^{2}}\Big)^{-1/2}\,d\tau$ and putting the expression for $u_{x}$ from equation (11) we can re-write equation (12) as:
$dx$ $=c\,\tanh\left[\frac{r_{0}\omega}{c}\Big\{\cos(\omega\tau)\right]\frac{d\tau}{\sqrt{1-\frac{u_{x}^{2}}{c^{2}}}}$
$\text{or\,\,\,\,\,}x$ $=\int\frac{c\,\tanh\left[\frac{r_{0}\omega}{c}\Big\{\cos(\omega\tau)\right]}{\sqrt{1-\tanh^{2}\left[\frac{r_{0}\omega}{c}\cos(\omega\tau)\right]}}d\tau$ (13)
The right hand side (RHS) of the above equation (13) is analytically not integrable. However, as in this equation $r_{0}\omega$ represents the maximum value of the linear velocity of the frame executing simple harmonic motion (SHM), and for practical purposes this velocity will be much smaller than the speed of light $(c)$. Thus assuming $\frac{r_{0}\,\omega}{c}<<1$, we can write: $\frac{r_{0}\omega}{c}\cos(\omega\tau)\sim 0$ and hence the right hand side (RHS) of equation (13) can be expand in a series in the power of $r_{0}\omega/c$. Now expanding the integrand in the RHS of equation (13) and neglecting the terms higher than $(r_{0}\omega/c)^{3}$ in order, we can obtain from the above equation (13) as below:
$x$ $=c\,\int\Bigg[0+\frac{r_{0}\omega}{c}\cos(\omega\tau)+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\cdot 0+\frac{r_{0}^{3}\omega^{3}}{6c^{3}}\cos^{3}(\omega\tau)\Bigg]d\tau$
$=r_{0}\sin(\omega\tau)+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Bigg\{9\sin(\omega\tau)+\sin(3\omega\tau)\Bigg\}+K_{2}$ (14)
where $K_{2}$ is a constant of integration.
Additionally, since the relation between the proper time ($\tau$) and the coordinate time $(t)$ can be written as: $dt=\Big\{1-\frac{u_{x}^{2}}{c^{2}}\Big\}^{-1/2}\,d\tau$, hence using equation (11) we can calculate the transformation relation for the time $t$ in case of SHM in $X$-direction as follows:
$\int dt=\int\Big\{1-\frac{u_{x}^{2}}{c^{2}}\Big\}^{-1/2}\,d\tau$
or $t=\int\left[1-\tanh^{2}\left[\frac{r_{0}\omega}{c}\cos(\omega\tau)\right]\right]^{-1/2}d\tau$ (15)
Again, since as stated above, $r_{0}\omega/c<<1$ for all practical purposes, we have $\frac{r_{0}\omega}{c}\cos(\omega\tau)\sim 0$, and hence we can take an approximation by neglecting the terms higher than $(r_{0}\omega/c)^{3}$ as before. So as before, expanding the integrand in the RHS of equation (15) in Taylor series and utilizing the above approximation we can obtain from equation (15) as:
$t$ $\approx\int\left[1+\frac{r_{0}\omega}{c}\cdot 0+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\cos^{2}(\omega\tau)+\frac{r_{0}^{3}\omega^{3}}{6c^{3}}\cdot 0\right]d\tau$
$=\int\left[1+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\cos^{2}(\omega\tau)\right]d\tau$
$=\tau+\frac{r_{0}^{2}\omega}{8c^{2}}\Bigg\{2\omega\tau+\sin(2\omega\tau)\Bigg\}+K_{3}$ (16)
where $K_{3}$ is a constant of integration.
Now in the present context, as $\tau$ can be the time in the frame executing SHM, so we may replace $\tau$ by $t^{\prime}$, where $t^{\prime}$ represents the time as measured from the accelerated frame (as discussed just before equation (4)). Therefore, substituting $\tau=t^{\prime}$ we can obtain from equations (14) and (16) as follows:
$x$ $=r_{0}\sin(\omega t^{\prime})+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Bigg[9\sin(\omega t^{\prime})+\sin(3\omega t^{\prime})\Bigg]+K_{2}$ (17)
$\text{and\,\,\,\,\,\,}t$ $=t^{\prime}+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega t^{\prime}+\sin(2\omega t^{\prime})\Big\}+K_{3}$ (18)
Limiting conditions:
  1. (i)
    In the initial condition, if the angular frequency of the oscillating frame is equal to zero, i.e. if $\omega=0$ then the coordinates in both frames will coincide, i.e. $x=x^{\prime}$. So, using that condition, we obtain from equation (17) as: $K_{2}=x^{\prime}$.
  2. (ii)
    On the other hand, if there is no oscillation i.e. $\omega=0$ then the time in both frames $S$ and $S^{\prime}$ should be same, i.e. $t=t^{\prime}$. So, using that condition we can obtain the constant of integration from equation (18) as: $K_{3}=0$.
Therefore, putting $K_{2}=x^{\prime}$ and $K_{3}=0$ in equations (17) and (18) we obtain the coordinate transformation relation for a frame $S^{\prime}(t^{\prime},x^{\prime},y^{\prime},z^{\prime})$ undergoing simple harmonic motion (SHM) in the $X$-direction with respect to an inertial frame $S(t,x,y,z)$ as follows:
$t$ $=t^{\prime}+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega t^{\prime}+\sin(2\omega t^{\prime})\Big\}$ (19)
$x$ $=x^{\prime}+r_{0}\sin(\omega t^{\prime})+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\sin(\omega t^{\prime})+\sin(3\omega t^{\prime})\Big\}$ (20)
$y$ $=y^{\prime}$ (21)
$z$ $=z^{\prime}$ (22)
In the limit of $\omega\longrightarrow 0$ we obtain from equations (19) and (20):
$\lim_{\omega\to 0}t=t^{\prime}$ (23)
and $\lim_{\omega\to 0}x=x^{\prime}$ (24)
Equations (23) and (24) above confirm the limiting condition when there is no oscillation $(\omega=0)$.
Now, in order to find a relation between time periods in the two frames we proceed as follows. We consider that $T$ and $T^{\prime}$ are the time periods as observed from the frames $S$ and $S^{\prime}$ respectively. From equation (19) we can write as:
$dt$ $=\Big[1+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega+2\omega\cos(2\omega t^{\prime})\Big\}\Big]dt^{\prime}$ (25)
Then we integrate both side of equation (25), with limit 0 to $T$, for the unprimed and 0 to $T^{\prime}$, for the primed frame. i.e.
$\int_{0}^{T}dt$ $=\int_{0}^{T^{\prime}}\Big[1+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega+2\omega\cos(2\omega t^{\prime})\Big\}\Big]dt^{\prime}$
$\text{or\hskip 42.67912pt}T$ $=T^{\prime}+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega T^{\prime}+\sin(2\omega T^{\prime})\Big\}$ (26)
Now, we can clearly see that, $T^{\prime}$ is the time period as observed by an observer sitting in the accelerated frame itself. So, it can be identified as the same time period as the non-relativistic one. We denote this non-relativistic time period as $T_{0}$ and we can write $T^{\prime}=T_{0}$. Therefore, putting $T^{\prime}=T_{0}=2\pi/\omega$, we obtain from the above equation (26) as follows:
$\frac{T}{T_{0}}=1+\frac{1}{4}\frac{\omega^{2}}{c^{2}}r_{0}^{2}$ (27)
Equation (27) is almost identical to the approximate expressions obtained by other authors [13, 15, 18]:
$\frac{T}{T_{0}}\approx 1+\frac{3}{16}\frac{\omega^{2}}{c^{2}}r_{0}^{2}$ (28)
Comparing (27) and (28), one can say that, we obtained $(1/4)$ as the the co-effecient of the second term in RHS, corresponding to $(3/16)$ for the same quantity as obtained by other authors [13, 15, 18].
Thus, calculations done by us are in good agreement with those approximate expressions reported by other authors in the past.
2.2 The frame $S^{\prime}$ undergoing SHM in $Y$-direction with initial phase of $\pi/2$ radians:
The circular motion of an object orbiting on the $XY$ plane is best described by two simultaneous SHMs (simple harmonic motions) of same amplitudes in $X$ and $Y$ directions separated by a phase angle of $\pi/2$ radians. As we want the $X$ and $Y$ components together to represent a circular motion (with radius $r_{0}$), we consider a displacement $y^{\prime}=r_{0}\cos(\omega\tau)$ of oscillation in $Y$-direction that is separated by a phase angle of $\pi/2$ radians from $x^{\prime}$ such that the velocity $u^{\prime}_{y}=-r_{0}\omega\sin(\omega\tau)$ and the acceleration $a^{\prime}_{y}=-r_{0}\omega^{2}\cos(\omega\tau)$ .
Therefore, using a similar approach to that in the case of SHM in $X$-direction, we can also obtain the transformation relation of the time and the spatial coordinates for SHM in $Y$-direction. As in the present case, the motion is along the $Y$-direction, therefore coordinate transformation relations can be calculated by simply replacing $a_{x}$, $u_{x}$ and $a_{x}^{\prime}$ respectively with $a_{y}=\frac{du_{y}}{dt}$, $u_{y}$ and $a^{\prime}_{y}=-r_{0}\omega^{2}\cos(\omega\tau)$ in equation (4), where $u_{y}$ is the $Y$- component of the three velocity. So, we can write as earlier for SHM in $Y$-direction as follows:
$\frac{du_{y}}{\left(1-\frac{u_{y}^{2}}{c^{2}}\right)}=-r_{0}\omega^{2}\cos(\omega\tau)\,d\tau$ (29)
Now, integrating both sides of the equation (29), we obtain :
$\int\frac{du_{y}}{\left(1-\frac{u_{y}^{2}}{c^{2}}\right)}=-r_{0}\omega^{2}\int\cos(\omega\tau)\,d\tau$
or $c\,\tanh^{-1}\left(\frac{u_{y}}{c}\right)=-r_{0}\omega\sin(\omega\tau)+K_{4}$
or $u_{y}=c\,\tanh\Bigg[-\frac{r_{0}\omega}{c}\sin(\omega\tau)+\frac{K_{4}}{c}\Bigg]$ (30)
where $K_{4}$ is an integration constant. Now, again as before, we consider that the displacement vector (of the $S^{\prime}$ frame) is purely oscillatory in nature, so the three velocity of the $S^{\prime}$ frame with respect to the inertial observer ($S$ frame) should be also oscillatory in nature with zero base line. Actually, one can show that the function $\tanh{\left[-\frac{r_{0}\omega}{c}\sin(\omega\tau)+\frac{K_{4}}{c}\right]}$ is oscillatory in nature with a baseline which depends upon our choice of the constant $K_{4}$. But if we set $K_{4}=0$, then the baseline also vanishes to zero. So we can safely assume $K_{4}$ to be equal to 0.
Therefore, now equation (30) can be written by putting $K_{4}=0$ as follows:
$u_{y}$ $=\frac{dy}{dt}=c\,\tanh\Bigg[-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Bigg]$ (31)
$\text{or\,\,\,\,\,\,\,}dy$ $=c\,\tanh\Bigg[-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Bigg]dt$ (32)
Then using the relation $dt=\Big(1-\frac{u_{y}^{2}}{c^{2}}\Big)^{-1/2}\,d\tau$ for SHM in $Y$-direction and putting the expression for $u_{y}$ from equation (31), we integrate equation (32) as follows:
$y$ $=\int c\,\tanh\Bigg[-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Bigg]\frac{d\tau}{\sqrt{1-\frac{u_{y}^{2}}{c^{2}}}}$
$=\int\frac{c\,\tanh\Bigg[-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Bigg]}{\sqrt{1-\tanh^{2}\Bigg[-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Bigg]}}d\tau$ (33)
The right hand side of equation (33) is not integrable. However, as stated before, assuming $r_{0}\omega/c<<1$, we can write $\frac{r_{0}\omega}{c}\sin(\omega\tau)\sim 0$ and expand the RHS of equation (33) in a series in the power of $(r_{0}\omega/c)$. So expanding the RHS of equation (33) and neglecting the terms higher than $(r_{0}\omega/c)^{3}$ we obtain as follows:
$y$ $\approx c\int\Big[0+\frac{r_{0}\omega}{c}\Big\{-\sin(\omega\tau)\Big\}+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\cdot 0+\frac{r_{0}^{3}\omega^{3}}{6c^{3}}\Big\{-\sin(\omega\tau)\Big\}^{3}\Big]d\tau$
$=r_{0}\cos(\omega\tau)+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Bigg[9\cos(\omega\tau)-\cos(3\omega\tau)\Bigg]+K_{5}$ (34)
where $K_{5}$ is a constant of integration.
Additionally, as calculated earlier, the transformation relation of time $t$ for SHM in $Y$-direction can be obtained by using the equation (31) as below:
$\int dt=\int\Big\{1-\frac{u_{y}^{2}}{c^{2}}\Big\}^{-1/2}\,d\tau$
or $t=\int\left[1-\tanh^{2}\left[\frac{r_{0}\omega}{c}\Big\{-\sin(\omega\tau)\Big\}\right]\right]^{-1/2}d\tau$ (35)
Now since $r_{0}\omega/c<<1$, we can take an approximation as before and therefore we can write from equation (35) as follows:
$t$ $\approx\int\left[1+\frac{r_{0}\omega}{c}\cdot 0+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\Big\{-\sin(\omega\tau)\Big\}^{2}+\frac{r_{0}^{3}\omega^{3}}{6c^{3}}\cdot 0\right]d\tau$
$\text{or\,\,\,\, }t$ $\approx\tau+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega\tau-\sin(2\omega\tau)\Big\}+K_{6}$ (36)
where $K_{6}$ is a constant of integration.
Then in a similar way, since $\tau$ can be the time in the frame executing SHM, we can replace $\tau$ by $t^{\prime}$ (like was done before for the $X$-component), where $t^{\prime}$ represents the time measured from the accelerated frame. So, substituting $\tau=t^{\prime}$ we obtain from equations (34) and (36) as follows:
$y$ $=r_{0}\cos(\omega t^{\prime})+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Bigg[9\cos(\omega t^{\prime})-\cos(3\omega t^{\prime})\Bigg]+K_{5}$ (37)
$t$ $=t^{\prime}+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega t^{\prime}-\sin(2\omega t^{\prime})\Big\}+K_{6}$ (38)
Limiting conditions:
  1. (i)
    In the initial condition, since the SHM in the present case is a cosine function of $t^{\prime}$ and $\omega$, so if the angular frequency of the oscillating frame is equal to zero, then the coordinates in both frames will be related as: $y=r_{0}+y^{\prime}$. So, using that condition, we obtain the integration constant from equation (37) as: $K_{5}=y^{\prime}$.
  2. (ii)
    On the other hand, if there is no oscillation, i.e. $\omega=0$, then the times in both frames $S$ and $S^{\prime}$ should be the same, i.e. $t=t^{\prime}$. So, using that condition, we can obtain the constant of integration from equation (38) as: $K_{6}=0$.
Therefore, putting $K_{5}=y^{\prime}$ and $K_{6}=0$ in equations (37) and (38) we obtain the coordinate transformation relation for a frame $S^{\prime}(t^{\prime},x^{\prime},y^{\prime},z^{\prime})$ undergoing simple harmonic motion (SHM) in $Y$-direction with respect to an inertial frame $S(t,x,y,z)$ as follows:
$t$ $=t^{\prime}+\frac{r_{0}^{2}\omega}{8c^{2}}\Big\{2\omega t^{\prime}-\sin(2\omega t^{\prime})\Big\}$ (39)
$x$ $=x^{\prime}$ (40)
$y$ $=y^{\prime}+r_{0}\cos(\omega t^{\prime})+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\cos(\omega t^{\prime})-\cos(3\omega t^{\prime})\Big\}$ (41)
$z$ $=z^{\prime}$ (42)
In the limiting condition of $\omega\longrightarrow 0$ the equations (39) and (41) become:
$\lim_{\omega\to 0}t=t^{\prime}_{y}$ (43)
and $\lim_{\omega\to 0}y=y^{\prime}+r_{0}$ (44)
The above equations (43) and (44) confirm the limiting condition when there is no oscillation (i.e. $\omega=0$).
3 Transformation relations for the frame moving in a circular path
As stated above, our objective is to obtain a transformation relation between an inertial frame and a frame moving in a circular path on the $XY$ plane. So we now consider that the frame $S^{\prime}$ is moving in a circular path with radius $r_{0}$ where the respective coordinates of two frames are parallel to each other as shown in Fig. 1.
Figure 1: Schematic of $S^{\prime}$ frame moving in a circular path on the $XY$ plane.
Since the circular motion of an object orbiting on the $XY$ plane can be described by two simultaneous SHMs of the same amplitudes in the $X$ and $Y$ directions separated by a phase angle of $\pi/2$ radians, so the transformation relations of the $X$ and $Y$ coordinates between the inertial frame and the rotating frame can be obtained from equations (20) and (41); while the $Z$ coordinate will remain unaffected. However, in order to obtain the transformation of the time coordinate in the present case, we used the standard relation between the time coordinate $(t)$ and the proper time $(\tau)$ as follows:
$t=\int\Big[1-\frac{u^{2}}{c^{2}}\Big]^{-1/2}d\tau$ (45)
where $u$ is the velocity on the $XY$ plane in the $S$ frame that can be written in terms of its mutually perpendicular components using equations (11) and (31) as below:
$u^{2}$ $=u^{2}_{x}+u^{2}_{y}$
$=c^{2}\tanh^{2}\Big\{\frac{r_{0}\omega}{c}\cos(\omega\tau)\Big\}+c^{2}\tanh^{2}\Big\{-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Big\}$ (46)
Therefore, using equation (46) into equation (45) we obtain as follows:
$t$ $=\int\Big[1-\Big\{\tanh^{2}\Big\{\frac{r_{0}\omega}{c}\cos(\omega\tau)\Big\}+\tanh^{2}\Big\{-\frac{r_{0}\omega}{c}\sin(\omega\tau)\Big\}\Big\}\Big]^{-1/2}d\tau$ (47)
Then, as mentioned before, assuming $r_{0}\omega/c<<1$ we expand the RHS of equation (47) in a series in the power of $r_{0}\omega/c$ and neglecting the terms higher than $(r_{0}\omega/c)^{3}$ we obtain as follows:
$t$ $\approx\int\Big[1+\frac{r_{0}\omega}{c}\cdot 0+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\cdot\left(\sin^{2}(\omega\tau)+\cos^{2}(\omega\tau)\right)+\frac{r_{0}^{3}\omega^{3}}{6c^{3}}\cdot 0\Big]d\tau$
$=\int\Bigg[1+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\Bigg]d\tau$
$=\tau+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\cdot\tau+K_{7}$ (48)
where $K_{7}$ is an integration constant. However, as before, the proper time $\tau$ in the present case can be the time $(t^{\prime})$ in the frame ($S^{\prime}$) that is moving in a circular path, and hence the transformation of time (48) can be written as:
$t=t^{\prime}\Bigg\{1+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\Bigg\}+K_{7}$ (49)
On the other hand, in the initial condition, if there is no oscillation, i.e., if $\omega=0$ then $t=t^{\prime}$ and hence the integration constant $K_{7}$ must be zero. Therefore, the coordinate transformation relations for a frame $S^{\prime}$ orbiting in a circular path can be written from equations (20), (41) and (49) (by putting $K_{7}=0$) as follows:
$t$ $=t^{\prime}\left\{1+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\right\}$ (50)
$x$ $=x^{\prime}+r_{0}\sin(\omega t^{\prime})+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\sin(\omega t^{\prime})+\sin(3\omega t^{\prime})\Big\}$ (51)
$y$ $=y^{\prime}+r_{0}\cos(\omega t^{\prime})+\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\cos(\omega t^{\prime})-\cos(3\omega t^{\prime})\Big\}$ (52)
$\text{and\,\,\,\,\,}z$ $=z^{\prime}$ (53)
Further, we can also write the inverse coordinate transformation relations between an inertial frame and a frame moving in a circular path using equations (50)—(53) by putting $t^{\prime}=\lambda t$, where $\lambda=\left\{1+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\right\}^{-1}$ as follows:
$t^{\prime}$ $=t\left\{1+\frac{r_{0}^{2}\omega^{2}}{2c^{2}}\right\}^{-1}=t\,\lambda$ (54)
$x^{\prime}$ $=x-r_{0}\sin(\lambda\omega t)-\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\sin(\lambda\omega t)+\sin(3\lambda\omega t)\Big\}$ (55)
$y^{\prime}$ $=y-r_{0}\cos(\lambda\omega t)-\frac{r_{0}^{3}\omega^{2}}{72c^{2}}\Big\{9\cos(\lambda\omega t)-\cos(3\lambda\omega t)\Big\}$ (56)
$\text{and\,\,\,\,\,}z^{\prime}$ $=z$ (57)
4 Comparison with some existing relativistic coordinate transformations
As mentioned above, in relativity the coordinate transformation relations are important tools that enable us to understand the dynamics of the test particles or an event that occurred in a frame of reference having different states of motion. In literature there exists a good number of works related to the coordinate transformation for non-inertial frames. In 1987, Nelson [2, 3] obtained a general coordinate transformation for non-inertial frames having an arbitrary time dependent translational acceleration and angular velocity by generalizing the Lorentz transformation. Ghosh and Sen [19] discussed about the rotation of polarization vector of an electromagnetic wave and redshift for light emitted from (or received by) a rotating frame, utilizing the transformation equations as already available in Nelson [2, 3], Mashhoon [20] etc.
In the present work, we investigated the coordinate transformation relations between an inertial frame and a non-inertial frame having non-uniform acceleration, more specifically the frames executing SHM and UCM. To our knowledge, particularly on the transformation relations for a frame executing SHM, only a few works have been reported in literature. For example, Perepelkin et al. [16] discussed about the transformation relation when the velocity of the moving frame is a sinusoidal function of time. Here, the authors obtained a Lorentz-like general coordinate transformation relations for non-inertial frames (see equation (20) and (24) in Perepelkin et al.[16]) by using the generalized Galilean transformation. Then the authors considered a time dependent velocity and directly incorporated it into the generalized Lorentz-like transformation.
On the other hand, in the case of the transformation relation between an inertial frame and a frame moving in a circular path, Nikolic [21] used Nelson’s [2, 3] relations to obtain a transformation relation for a rotating frame which is similar to the transformation relations obtained by Mashhoon and Muench [22].
Additionally, Kipreos and Balachandran [23] compared some of the then existing relativistic coordinate transformation relations for rotating frames available in the literature. Here, the authors primarily considered four transformation relations, namely, Post transformation, Franklin transformation, the rotational form of the absolute Lorentz transformation (ALT) and the Langevin metric. And then they compared the theoretical predictions given by those transformation relations with the data including the observations of length contraction, directional time dilation, anisotropic one-way speed of light, isotropic two-way speed of light, and the conventional Sagnac effect. Most of the transformation relations that the authors [23] compared were based on certain assumptions. However, in our present work we believe we extended Lorentz/ Rindler’s transformation equations in a systematic manner and obtained certain analytical expressions. So we think our work can not be directly compared with the set of various similar work reported in Kipreos and Balachandran[23]. The relativistic transformation for a rotating frame obtained by using ALT (following Franklin [24]), as has been reported in Kipreos and Balachandran[23] is probably the best one as it matched most with the observational data. Interestingly, the Langevin relativistic transformation in Cartesian coordinates can be written as follows [25]:
[…]
which is similar to our results shown in equations (50)—(53) if we neglect the higher order terms containing $r_{0}^{3}\omega^{2}/(72c^{2})$. However, in our coordinate transformation equations, for the frame executing UCM there exist higher order terms containing $r_{0}^{3}\omega^{2}/(72c^{2})$ which indicate the relativistic anharmonic effect on the circular motion.
5 Discussion of results and conclusions
In this work we obtained the transformation relations between an inertial frame and a frame executing non-uniform acceleration such as SHM. In a similar way, calculations were also done to determine the coordinate transformation relations between an inertial frame and another frame having uniform circular motion (UCM). These transformation relations can be used to investigate the dynamics of the test particles in the above mentioned non-uniformly accelerated frames.
Additionally, the results in this paper will be applicable for practical calculations of the relativistic effects, such as in a spacecraft orbiting around the Earth, etc. Apart from that, these calculations will have applications in different branches of Relativistic Physics and more specifically in Relativistic Astrophysics and General Relativity. For example, materials in the circumstellar shells of pulsating variable stars undergoing simple harmonic motions. In addition we have many binary stars with the components moving in circular orbits around the common center of mass. When we study their dynamics, or dynamics of an electron or photon (light) in the gravitational fields of such binary stars, then the equations derived in our present work will be very useful.
Acknowledgments
The authors thank faculty members and colleagues of the Department of Physics, Assam University, Silchar-788011 (India) for their encouragement and support while doing this work.